OpenStudy (callisto):

Graph problem If we want to shift the graph to the left, we would write the equation y=f(x) into y=f(x+k), where k is a positive value. What about if I want to shift the graph of y=e^[f(x)] by k units to the left. What should I do? Should I do it in the normal way?

OpenStudy (inkyvoyd):

y=e^[f(x)]?

OpenStudy (callisto):

Yes.. If I want to shift this graph by k units to the left

OpenStudy (inkyvoyd):

What's f(x)? Just a general function?

OpenStudy (callisto):

Yes

OpenStudy (inkyvoyd):

Well, same thing.

OpenStudy (callisto):

Really?!

OpenStudy (inkyvoyd):

Wait.

OpenStudy (inkyvoyd):

say f(x)=x

OpenStudy (inkyvoyd):

Let me test this out with geogebra.

OpenStudy (inkyvoyd):

Yup.

OpenStudy (inkyvoyd):

Assuming f(x) is a function of x, ou basically have f(x+k)

OpenStudy (callisto):

So the graph would be y=e^[f(x+k)]?

OpenStudy (inkyvoyd):

Yes

OpenStudy (inkyvoyd):

if f(x)=e^x, you have e^(e^(x+k))

OpenStudy (callisto):

Wow~ that's weird!!! so, y = f(x) = e^[f(x)] = f(x+k) = e^ (e^(x+k)) ??

OpenStudy (inkyvoyd):

Think about it. For any function f(x), if you add a number k to every x, then f(x) needs k less input to get output f1(x).

OpenStudy (inkyvoyd):

If you subtract k from every input x, however, then f(x) needs k more input to get output f2(x). thus for the former case the graph is shifted left, and for the latter case the graph is shifted right, because it requires less to be larger (in the case of e^x, all other functions work, monotonic or not) and because it requires more to be larger (also in the case of e^x), respectively.

OpenStudy (inkyvoyd):

If you have a polynomial, say x^3-x^2+2x=f(x), f(x+k) would be (x+3)^3-(x+3)^2+2(x+3)

OpenStudy (inkyvoyd):



OpenStudy (callisto):

I think I understand the general case. Was not sure when the power is the function one..

OpenStudy (inkyvoyd):